OpenStudy (anonymous):

Solve the system of equations. 3x - 4y = -10 and 5x + 8y = -2

hero (hero):

3x - 4y = -10 5x + 8y = -2 How come you don't want to learn how to solve these yourself?

OpenStudy (anonymous):

Well I WOULD like to, but I'm kind of in a hurry right now.

hero (hero):

You always seem to be in a hurry

OpenStudy (anonymous):

Well, I'm actually really in a hurry this time.

OpenStudy (anonymous):

3x - 4y = -10 (Eqn. 1) 5x + 8y = -2 (Eqn. 2) Right. Multiply equation 1 by 2 to get: 6x - 8y = -20 5x + 8y = -2 (Eqn. 2 stays the same) Now, 8 appears in both equations but they have different signs so add equations 1 and 2 to get: 11x + 0y = -22 11x = -22 x = -2 Sub x = -2 into one of the original equations to find y. We'll use Eqn. 2. So: 5x + 8y = -2 5(-2) + 8y = -2 -10 + 8y = -2 8y = 8 y = 1 There we go. :)

OpenStudy (anonymous):

That's the answer...